Solve by gaussian method

X+2Y-3Y=11

3X+2Y+Z=1

2X+Y-5Z=11


1
Expert's answer
2021-06-11T03:02:55-0400

Augmented matrix


"\begin{pmatrix}
 1 & 2 & -3 & & 11 \\
 3 & 2 & 1 & & 1 \\
 2 & 1 & -5 & & 11 \\
\end{pmatrix}"



"R_2=R_2-3R_1"


"\begin{pmatrix}
 1 & 2 & -3 & & 11 \\
 0 & -4 & 10 & & -32 \\
 2 & 1 & -5 & & 11 \\
\end{pmatrix}"


"R_3=R_3-2R_1"


"\begin{pmatrix}
 1 & 2 & -3 & & 11 \\
 0 & -4 & 10 & & -32 \\
 0 & -3 & 1 & & -11 \\
\end{pmatrix}"



"R_2=-R_2\/4"


"\begin{pmatrix}
 1 & 2 & -3 & & 11 \\
 0 & 1 & -5/2 & & 8 \\
 0 & -3 & 1 & & -11 \\
\end{pmatrix}"

"R_1=R_1-2R_2"


"\begin{pmatrix}
 1 & 0 & 2 & & -5 \\
 0 & 1 & -5/2 & & 8 \\
 0 & -3 & 1 & & -11 \\
\end{pmatrix}"

"R_3=R_3+3R_2"


"\begin{pmatrix}
 1 & 0 & 2 & & -5 \\
 0 & 1 & -5/2 & & 8 \\
 0 & 0 & -13/2 & & 13 \\
\end{pmatrix}"

"R_3=-(2\/13)R_3"


"\begin{pmatrix}
 1 & 0 & 2 & & -5 \\
 0 & 1 & -5/2 & & 8 \\
 0 & 0 & 1 & & -2\\
\end{pmatrix}"

"R_1=R_1-2R_3"


"\begin{pmatrix}
 1 & 0 & 0 & & -1 \\
 0 & 1 & -5/2 & & 8 \\
 0 & 0 & 1 & & -2\\
\end{pmatrix}"

"R_2=R_2+(5\/2)R_3"


"\begin{pmatrix}
 1 & 0 & 0 & & -1 \\
 0 & 1 & 0 & & 3 \\
 0 & 0 & 1 & & -2\\
\end{pmatrix}"

Then


"x=-1, y=3, z=-2"

"(-1, 3, -2)"